A sound source of frequency 1000 Hz moves at 40.0 m/s toward a listener who is at rest. What is the apparent frequency heard by the listener? (speed of sound = 340 m/s)


1
Expert's answer
2022-01-20T13:52:35-0500

Answer


The Doppler's effect,

when source is moving towards the at rest (stationary) observer, then observed frequency is given by:

"f = f0*\frac{(V - Vo)}{(V - Vs)}"

V = Speed of sound = 340 m/s

Vo = speed of observer = 0

Vs = speed of source = 40.0 m/s

f0 = frequency of source = 1000 Hz

So,

Putting all value

"f = 1100*\frac{(340 - 0)}{(340 - 40.0)}\\

f = 1246.7 Hz\\ = 1250 Hz"